Write the Lagrangian of this particle. Show that the Euler-Lagrange equa-

tions of motion for this particle is identical to what one would obtain from

Newton’s second law.


1
Expert's answer
2021-08-16T08:37:09-0400

The lagrangian of the system

"L=T-V=\frac{m({\dot x}^2+{\dot y}^2)}{2}-mgy"

The Euler-Lagrange equations

"\frac{d}{dt}\left(\frac{\partial L}{\partial {\dot x}}\right)-\frac{\partial L}{\partial {x}}=0,\\\frac{d}{dt}\left(\frac{\partial L}{\partial {\dot y}}\right)-\frac{\partial L}{\partial {y}}=0,"

give


"m{\ddot x}+0=0,\\
m{\ddot y}+mg=0."

Finally


"m{\ddot x}=0,\\
m{\ddot y}=-mg."